# Minimal ring extensions of the integers exhibiting Kochen-Specker contextuality

### v3

@@ 2.2. A computational method for detecting uncolorability
Our further explorations into the uncolorability of $\mathcal{S}_{d}(N)$ were aided by computer searches. This consists of generating the vectors in $v\in\mathcal{S}_{d}(N)$ such that $q(v)$ divides $N^{e}$ for some power $e\geq 1$, and testing whether the set of vectors is colorable or not. In order to test colorability of the vectors, we have found it useful to apply the method of [20, Construction 3.1.3] to convert the question of whether a Kochen-Specker coloring exists into an integer linear programming problem. We describe that method below.
@@
We implemented this method using Python code. [^2] The code is publicly available at https://github.com/manny-reyes/Kochen_Specker_Colorability. For a positive integer $N$ and dimension $3\leq d\leq 6$, the algorithm proceeds as follows:
@@
Using this code we have been able to explore whether $\mathcal{S}_{d}(N)$ is colorable for various values of $N$ across dimensions $3\leq d\leq 6$. Some of our computational findings in dimension $d=3$ are summarized in Table 1. Our computations verify those of [20, §3.3.1] for all overlapping values of $N$, including the interesting fact that $\mathcal{S}_{3}(714)$ is uncolorable.
@@ 2.3. Uncolorability in dimensions $d\geq 4$
The method of proof is similar to that of [5, Lemma 2.17]. Assume that $\mathcal{S}_{d}(N)$ is KS-uncolorable; proceeding by induction, it suffices to show that $\mathcal{S}_{d+1}(N)$ is also KS-uncolorable.
@@
We have used the method of Subsection 2.2 to investigate colorability of $\mathcal{S}_{d}(N)$ for the range of dimensions $d=4,5,6$. Some of these findings are summarized in Table 2. The most notable discovery is the uncolorability of $\mathcal{S}_{6}(3)$, which we record in the theorem below. The uncolorable set consists of those vectors $v\in\mathbb{Z}^{6}$ such that $q(v)\in\{1,3\}$. Choosing a single representative out of each pair ±$\pm v$ of such vectors on the same line yields a set of 86 vectors. (This includes the six standard basis vectors, and $\binom{6}{3}\cdot 2^{2}=80$ other vectors with three nonzero entries that are ±$\pm 1$ with, say, their first entries positive.) We have not searched to find the smallest possible uncolorable subset of these vectors.
@@ 6 $p\in\{5,7,11,\dots,19\}$ 1 colorable
(1) As noted earlier, there is a KS uncolorable set $S\subseteq\mathbb{Z}^{4}$ constructed in [16] such that $N(S)=4$. Thus $\mathcal{S}_{4}(2)\supseteq S$ is also uncolorable. The case $d\geq 4$ now follows from Lemma 4.
(2) As shown in Table 2, the set of vectors $v\in\mathcal{S}_{3}(6)$ satisfying $q(v)\in\{1,3\}$ is KS uncolorable. So the result follows from Lemma 4 again. ∎
The dimension $d=4,5$ results in Table 2 suggested to us the possibility that $\mathcal{S}_{d}(N)$ might be KS-colorable for odd $N$ in these dimensions. This led us to apply the method described Subsection 2.2 to investigate whether the projection matrices over the field $\mathbb{F}_{2}$ with two elements were Kochen-Specker colorable. [^3] The code is publicly available at https://github.com/manny-reyes/Kochen_Specker_Colorability. Our findings are recorded in Table 3.
Table 3. Colorability of $\operatorname{Proj}(\mathbb{M}_{d}(\mathbb{F}_{2}))$
@@
$\mathcal{S}_{d}(\mathbb{F}_{2})$ =$=\{v\in\mathbb{F}_{2}^{d}\mid v\cdot v\neq 0\}$
[…]
@@ 3. Consequences for symmetric matrices over commutative rings
We now apply the results of Section 2 to the study of contextuality in the purely algebraic setting, culminating in a proof of Theorem 1. The link from vector colorings to algebraic hidden states occurs in a multi-step process: uncolorability of vectors implies uncolorability of projection matrices, which in turn implies that a certain spectrum is empty (Lemma 8), which finally implies the non-existence of morphisms to commutative rings. To explain these connnections, we briefly provide several definitions from [10, Section 2] and [5, Section 2], to which readers are referred for more details.
A partial Boolean algebra is a set $B$ equipped with a reflexive, symmetric binary operation $\odot$ of commesaurability, distinguished elements $0,1\in B$, a unary operation $\neg$ of negation, and partially defined binary operations of meet $\wedge$ and join $\vee$ for commeasurable pairs of elements, such that every pairwise commeasurable set $S\subseteq B$ is contained in a pairwise commeasurable set $C\subseteq B$ containing 00 and $1$ for which the restricted negation, meet, and join make $C$ into an ordinary (“total”) Boolean algebra. The colorings of Kochen and Specker correspond to homomorphisms of partial Boolean algebras $B\to\mathbf{2}$, where $\mathbf{2}=\{0,1\}$ is the ordinary two-element Boolean algebra. We simply refer to such morphisms as Kochen-Specker colorings of a partial Boolean algebra as in [5, Theorem 2.13].
@@
Bridging the gap between colorability of projections and morphisms of partial rings requires prime partial ideals, as defined in [18]. Let $R$ be a partial ring. A subset $I\subseteq R$ is a partial ideal if it satisfies the following conditions for all $a,b\in R$:
@@
The partial spectrum $\operatorname{\mathit{p}-Spec}(R)$ is defined to be the set of all prime partial ideals of $R$. This forms a contravariant functor from the category of partial rings to the category of sets in the usual way, where a morphism $f\colon R\to S$ induces a map $\operatorname{\mathit{p}-Spec}(S)\to\operatorname{\mathit{p}-Spec}(R)$ by $P\mapsto f^{-1}(P)$. The link between the partial spectrum of $R$ and Kochen-Specker colorings is given by the following observation from [5, (2.15)] . If there exists a prime partial ideal $P\in\operatorname{\mathit{p}-Spec}(R)$, then we obtain a Kochen-Specker coloring on the partial Boolean algebra $B=\operatorname{Idem}(R)$ of idempotents in $R$, given by the following assignment for $e\in B$:
@@
On the other hand, if $\operatorname{\mathit{p}-Spec}(R)=\varnothing$ then there is no morphism of partial rings $R\to C$ for any commutative (total) ring $C$ by [5, Lemma 2.4]. In this way, the non-existence of prime partial ideals can be viewed as an algebraic manifestation of contextuality.
Note that if $K$ is a commutative ring and $R=\mathbb{M}_{n}(K)_{\mathrm{sym}}$ is a partial algebra of symmetric matrices, then the partial Boolean algebra of idempotents recovers the partial Boolean algebra of projections discussed above: $\operatorname{Idem}(R)=\operatorname{Proj}(\mathbb{M}_{n}(K))$. Thus uncolorability of $\operatorname{Proj}(\mathbb{M}_{n}(K))$ implies that there is no morphism of partial rings $\mathbb{M}_{n}(K)_{\mathrm{sym}}\to C$ for any commutative ring $C$.
@@
(i)$\implies$(ii): Each vector in $\mathcal{S}_{d}(N)$ defines a rank-1 projection in $\mathbb{M}_{d}(\mathbb{Z}[1/N])$, with orthogonality of vectors corresponds to orthogonality of the associated projections. Thus every Kochen-Specker coloring $\operatorname{Proj}(\mathbb{M}_{d}(\mathbb{Z}[1/N]))\to\{0,1\}$ induces a corresponding KS coloring of $\mathcal{S}_{d}(N)$. If $\mathcal{S}_{d}(N)$ is KS uncolorable, this means that the projections of $\mathbb{M}_{d}(\mathbb{Z}[1/N])$ are also KS uncolorable.
(ii)$\implies$(iii) is proved in [5, Corollary 2.16]. ∎
@@
(1) If $N$ is not divisible by $6$, then it is relatively prime to 2 or 3 (possibly both). This means that $N$ is a unit modulo 2 or 3, so that we have a ring homomorphism from $\mathbb{Z}[1/N]\to\mathbb{F}_{p}$, and consequently a morphism $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}\to\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}$, for either $p=2,3$. By [5, Theorem 3.4] there exists a prime partial ideal of $\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}$ for each of these values of $p$, and its preimage gives a prime partial ideal of $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}$.
Now suppose that $N$ is divisible by 6. By [6], the set $\mathcal{S}_{3}(30)$ has no Kochen-Specker coloring, and by Theorem 3 the set $\mathcal{S}_{3}(462)$ has no Kochen-Specker coloring. Because $6=\gcd(30,462)$, it follows from Lemma 9 that $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}$ has no prime partial ideals.
@@
(3) Uncolorability of $\mathcal{S}_{4}(2)$ as in part (2) above implies that $\operatorname{Proj}(\mathbb{M}_{4}(\mathbb{Z}[1/2]))$ has no KS coloring. It follows from [5, Lemma 2.17(3)] that $\operatorname{Proj}(\mathbb{M}_{6}(Z[1/2])$ also has no KS coloring. In particular, $\mathcal{S}_{6}(2)$ has no KS coloring. As shown in Theorem 5, $\mathcal{S}_{6}(3)$ also has no KS coloring. Because $1=\gcd(2,3)$, Lemmas 8 and 9 imply that $\mathbb{M}_{6}(\mathbb{Z})_{\mathrm{sym}}$ has no prime partial ideals. The claim for $\mathbb{M}_{d}(\mathbb{Z})_{\mathrm{sym}}$ with $d\geq 6$ now follows from [5, Lemma 2.17(4)]. ∎
@@
(1) First suppose that $6\mid N$. By Theorem 10 and [5, Lemma 2.4], there is no morphism of partial rings $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}\to C$ for any commutative ring $C$.
Now suppose that $p\notdivides N$ for either $p=2,3$. Then the (unique) ring homomorphism $\mathbb{Z}[1/N]\to\mathbb{F}_{p}$ induces a morphism of partial rings $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}\to\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}$. There exists a morphism of partial rings $\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}\to\mathbb{F}_{p^{6}}$ as in the proof of [5, Theorem 3.5]. Composing these maps gives a morphism of partial rings
@@
(2–3) The nonexistence of morphisms of partial rings in these cases follows directly from [5, Lemma 2.4] and Theorem 10. ∎
In Theorem 1(2), there is the remaining question of whether one can construct an algebraic hidden state for $\mathbb{M}_{d}(\mathbb{Z}[1/2])_{\mathrm{sym}}$ in dimensions $d=4,5$. The construction in the case of dimension 3 began with a coloring $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{F}_{3}))$, which was extended using [5, Appendix A] to a morphism of partial rings to an extension of $\mathbb{F}_{3}$. But that method of extension relied crucially on the 3-dimensionality of the situation. It is conceivable that a similar method could leverage a coloring such as that of Proposition 6, but it is . But such an approach would still require some fresh insights, so we do not pursue it further in this paper.
The results of Section 2 also allow us to investigate KS colorability of projections over some rings of number-theoretic interest. This was explored to some extent in [5], including for finite fields of prime order. We are now able to extend this treatment to $p$-adic numbers and localizations of the ring of integers in a unified way. For a prime $p$, let $\mathbb{F}_{p}=\mathbb{Z}/p\mathbb{Z}$ denote the field with $p$ elements, let $\mathbb{Z}_{p}$ denote the ring of $p$-adic integers, and let $\mathbb{Z}_{(p)}=\{a/b\in\mathbb{Q}\mid a,b\in\mathbb{Z},\,p\notdivides b\}$ denote the localization of the integers at the prime ideal $(p)$.
@@
For $p=2,3$ it follows from [5, Theorem 3.4] that $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{F}_{p}))$ has a Kochen-Specker coloring. By lifting along the morphism $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{Z}_{p}))\to\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{F}_{p}))$ (see also [5, Theorem 2.13]), we obtain a coloring of the $3\times 3$ projection matrices over $\mathbb{Z}_{p}$ and $\mathbb{Z}_{(p)}$.
For $p=5$, we have $\mathbb{Z}[1/462]\subseteq\mathbb{Z}_{(5)}$ and for $p>5$ we have $\mathbb{Z}[1/30]\subseteq\mathbb{Z}_{(p)}$. In light of Lemma 8, the $3\times 3$ projection matrices over $\mathbb{Z}[1/462]$ are uncolorable by Theorem 3, while those over $\mathbb{Z}[1/30]$ are uncolorable by [6]. In either case, we can embed an uncolorable set in $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{Z}_{(p)}))$, and the image of this set remains uncolorable over $\mathbb{Z}_{p}$ and $\mathbb{F}_{p}$.
@@
However, our computational exploration (Table 1) has failed to produce an uncolorable set in $\mathcal{S}_{3}(6)$ to very large norm, suggesting that perhaps this set has a KS coloring. If $\mathcal{S}_{3}(6)$ turns out to be KS colorable, it is far from clear what property would separate those values of $N$ that produce KS colorable sets from those that do not. For instance, it was shown by Salt [20, Table 3.2] that $\mathcal{S}_{3}(714)$ is also KS uncolorable, where $714=2\cdot 3\cdot 7\cdot 17$.

### v4

@@ 2.2. A computational method for detecting uncolorability
Our further explorations into the uncolorability of $\mathcal{S}_{d}(N)$ were aided by computer searches. This consists of generating the vectors $v\in\mathcal{S}_{d}(N)$ such that $q(v)$ divides $N^{e}$ for some power $e\geq 1$, and testing whether the set of vectors is colorable or not. In order to test colorability of the vectors, we have found it useful to apply the method of [Salt, Construction 3.1.3] to convert the question of whether a Kochen-Specker coloring exists into an integer linear programming problem. We describe that method below.
@@
We implemented this method using Python code. [^2] The code can be found in the folder https://github.com/manny-reyes/Kochen_Specker_Colorability/tree/main/Dimension_d_vector_coloring. For a positive integer $N$ and dimension $3\leq d\leq 6$, the algorithm proceeds as follows:
@@
Using this code we have been able to explore whether $\mathcal{S}_{d}(N)$ is colorable for various values of $N$ across dimensions $3\leq d\leq 6$. Some of our computational findings in dimension $d=3$ are summarized in Table 1. Our computations verify those of [Salt, §3.3.1] for all overlapping values of $N$, including the interesting fact that $\mathcal{S}_{3}(714)$ is uncolorable.
@@ 2.3. Uncolorability in dimensions $d\geq 4$
The method of proof is similar to that of [BMR, Lemma 2.17]. Assume that $\mathcal{S}_{d}(N)$ is KS-uncolorable; proceeding by induction, it suffices to show that $\mathcal{S}_{d+1}(N)$ is also KS-uncolorable.
@@
We have used the method of Subsection 2.2 to investigate colorability of $\mathcal{S}_{d}(N)$ for the range of dimensions $d=4,5,6$. Some of these findings are summarized in Table 2. The most notable discovery is the uncolorability of $\mathcal{S}_{6}(3)$, which we record in the theorem below. The uncolorable set consists of those vectors $v\in\mathbb{Z}^{6}$ such that $q(v)\in\{1,3\}$. Choosing a single representative out of each pair $\pm v$ of such vectors on the same line yields a set of 86 vectors. (This includes the six standard basis vectors, and $\binom{6}{3}\cdot 2^{2}=80$ other vectors with three nonzero entries that are $\pm 1$ with, say, their first entries positive.) We have not searched to find the smallest possible uncolorable subset of these vectors.
@@ 6 $p\in\{5,7,11,\dots,19\}$ 1 colorable
(1) As noted earlier, there is a KS uncolorable set $S\subseteq\mathbb{Z}^{4}$ constructed in [Peres:KS] such that $N(S)=4$. Thus $\mathcal{S}_{4}(2)\supseteq S$ is also uncolorable. The case $d\geq 4$ now follows from Lemma 4.
(2) As shown in Table 2, the set of vectors $v\in\mathcal{S}_{6}(3)$ satisfying $q(v)\in\{1,3\}$ is KS uncolorable. So the result follows from Lemma 4 again. ∎
The dimension $d=4,5$ results in Table 2 suggested to us the possibility that $\mathcal{S}_{d}(N)$ might be KS-colorable for odd $N$ in these dimensions. This led us to apply the method described Subsection 2.2 to investigate whether the projection matrices over the field $\mathbb{F}_{2}$ with two elements were Kochen-Specker colorable. [^3] This code is available at https://github.com/manny-reyes/Kochen_Specker_Colorability/tree/main/Projection_coloring. Our findings are recorded in Table 3.
Table 3. Colorability of $\operatorname{Proj}(\mathbb{M}_{d}(\mathbb{F}_{2}))$.
@@
$\mathcal{S}_{d}(\mathbb{F}_{2})$ $=\{v\in\mathbb{F}_{2}^{d}\mid v\cdot v\neq 0\}$
$=\{v\in\mathbb{F}_{2}^{d}\mid v\mbox{ has an odd number of nonzero entries}\}.$
@@ 3. Consequences for symmetric matrices over commutative rings
We now apply the results of Section 2 to the study of contextuality in the purely algebraic setting, culminating in a proof of Theorem 1. The link from vector colorings to algebraic hidden states occurs in a multi-step process: uncolorability of vectors implies uncolorability of projection matrices, which in turn implies that a certain spectrum is empty (Lemma 8), which finally implies the non-existence of morphisms to commutative rings. To explain these connnections, we briefly provide several definitions from [KochenSpecker, Section 2] and [BMR, Section 2], to which readers are referred for more details.
A partial Boolean algebra is a set $B$ equipped with a reflexive, symmetric binary operation $\odot$ of commesaurability, distinguished elements $0,1\in B$, a unary operation $\neg$ of negation, and partially defined binary operations of meet $\wedge$ and join $\vee$ for commeasurable pairs of elements, such that every pairwise commeasurable set $S\subseteq B$ is contained in a pairwise commeasurable set $C\subseteq B$ containing 0 and $1$ for which the restricted negation, meet, and join make $C$ into an ordinary (“total”) Boolean algebra. The colorings of Kochen and Specker correspond to homomorphisms of partial Boolean algebras $B\to\mathbf{2}$, where $\mathbf{2}=\{0,1\}$ is the ordinary two-element Boolean algebra. We simply refer to such morphisms as Kochen-Specker colorings of a partial Boolean algebra as in [BMR, Theorem 2.13].
@@
Bridging the gap between colorability of projections and morphisms of partial rings requires prime partial ideals, as defined in [Reyes]. Let $R$ be a partial ring. A subset $I\subseteq R$ is a partial ideal if it satisfies the following conditions for all $a,b\in R$:
@@
The partial spectrum $\operatorname{\mathit{p}-Spec}(R)$ is defined to be the set of all prime partial ideals of $R$. This forms a contravariant functor from the category of partial rings to the category of sets in the usual way, where a morphism $f\colon R\to S$ induces a map $\operatorname{\mathit{p}-Spec}(S)\to\operatorname{\mathit{p}-Spec}(R)$ by $P\mapsto f^{-1}(P)$. The link between the partial spectrum of $R$ and Kochen-Specker colorings is given by the following observation from [BMR, (2.15)] . If there exists a prime partial ideal $P\in\operatorname{\mathit{p}-Spec}(R)$, then we obtain a Kochen-Specker coloring on the partial Boolean algebra $B=\operatorname{Idem}(R)$ of idempotents in $R$, given by the following assignment for $e\in B$:
@@
On the other hand, if $\operatorname{\mathit{p}-Spec}(R)=\varnothing$ then there is no morphism of partial rings $R\to C$ for any nonzero commutative (total) ring $C$ by [BMR, Lemma 2.4]. In this way, the non-existence of prime partial ideals can be viewed as an algebraic manifestation of contextuality.
Note that if $K$ is a commutative ring and $R=\mathbb{M}_{n}(K)_{\mathrm{sym}}$ is a partial algebra of symmetric matrices, then the partial Boolean algebra of idempotents recovers the partial Boolean algebra of projections discussed above: $\operatorname{Idem}(R)=\operatorname{Proj}(\mathbb{M}_{n}(K))$. Thus uncolorability of $\operatorname{Proj}(\mathbb{M}_{n}(K))$ implies that there is no morphism of partial rings $\mathbb{M}_{n}(K)_{\mathrm{sym}}\to C$ for any nonzero commutative ring $C$.
@@
[…]
@@
(1) If $N$ is not divisible by $6$, then it is relatively prime to 2 or 3 (possibly both). This means that $N$ is a unit modulo 2 or 3, so that we have a ring homomorphism from $\mathbb{Z}[1/N]\to\mathbb{F}_{p}$, and consequently a morphism $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}\to\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}$, for either $p=2,3$. By [BMR, Theorem 3.4] there exists a prime partial ideal of $\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}$ for each of these values of $p$, and its preimage gives a prime partial ideal of $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}$.
Now suppose that $N$ is divisible by 6. By [Bub:Schutte], the set $\mathcal{S}_{3}(30)$ has no Kochen-Specker coloring, and by Theorem 3 the set $\mathcal{S}_{3}(462)$ has no Kochen-Specker coloring. Because $6=\gcd(30,462)$, it follows from Lemma 9 that $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}$ has no prime partial ideals.
@@
(3) Uncolorability of $\mathcal{S}_{4}(2)$ as in part (2) above implies that $\operatorname{Proj}(\mathbb{M}_{4}(\mathbb{Z}[1/2]))$ has no KS coloring. It follows from [BMR, Lemma 2.17(3)] that $\operatorname{Proj}(\mathbb{M}_{6}(Z[1/2])$ also has no KS coloring. As shown in Theorem 5, $\mathcal{S}_{6}(3)$ also has no KS coloring. Because $1=\gcd(2,3)$, Lemmas 8 and 9 imply that $\mathbb{M}_{6}(\mathbb{Z})_{\mathrm{sym}}$ has no prime partial ideals. The claim for $\mathbb{M}_{d}(\mathbb{Z})_{\mathrm{sym}}$ with $d\geq 6$ now follows from [BMR, Lemma 2.17(4)]. ∎
@@
(1) First suppose that $6\mid N$. By Theorem 10 and [BMR, Lemma 2.4], there is no morphism of partial rings $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}\to C$ for any nonzero commutative ring $C$.
Now suppose that $p\notdivides N$ for either $p=2,3$. Then the (unique) ring homomorphism $\mathbb{Z}[1/N]\to\mathbb{F}_{p}$ induces a morphism of partial rings $\mathbb{M}_{3}(\mathbb{Z}[1/N])_{\mathrm{sym}}\to\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}$. There exists a morphism of partial rings $\mathbb{M}_{3}(\mathbb{F}_{p})_{\mathrm{sym}}\to\mathbb{F}_{p^{6}}$ as in the proof of [BMR, Theorem 3.5]. Composing these maps gives a morphism of partial rings
@@
(2–3) The nonexistence of morphisms of partial rings in these cases follows directly from [BMR, Lemma 2.4] and Theorem 10. ∎
In Theorem 1(2), there is the remaining question of whether one can construct an algebraic hidden state for $\mathbb{M}_{d}(\mathbb{Z}[1/2])_{\mathrm{sym}}$ in dimensions $d=4,5$. The construction in the case of dimension 3 began with a coloring $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{F}_{3}))$, which was extended using [BMR, Appendix A] to a morphism of partial rings to an extension of $\mathbb{F}_{3}$. But that method of extension relied crucially on the 3-dimensionality of the situation. It is conceivable that a similar method could leverage a coloring such as that of Proposition 6 to define a morphism of partial rings. But such an approach would still require some fresh insights, so we do not pursue it further in this paper.
The results of Section 2 also allow us to investigate KS colorability of projections over some rings of number-theoretic interest. This was explored to some extent in [BMR], including for finite fields of prime order. We are now able to extend this treatment to $p$-adic numbers and localizations of the ring of integers in a unified way. For a prime $p$, let $\mathbb{F}_{p}=\mathbb{Z}/p\mathbb{Z}$ denote the field with $p$ elements, let $\mathbb{Z}_{p}$ denote the ring of $p$-adic integers, and let $\mathbb{Z}_{(p)}=\{a/b\in\mathbb{Q}\mid a,b\in\mathbb{Z},\,p\notdivides b\}$ denote the localization of the integers at the prime ideal $(p)$.
@@
For $p=2,3$ it follows from [BMR, Theorem 3.4] that $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{F}_{p}))$ has a Kochen-Specker coloring. By lifting along the morphism $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{Z}_{p}))\to\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{F}_{p}))$ (see also [BMR, Theorem 2.13]), we obtain a coloring of the $3\times 3$ projection matrices over $\mathbb{Z}_{p}$ and $\mathbb{Z}_{(p)}$.
For $p=5$, we have $\mathbb{Z}[1/462]\subseteq\mathbb{Z}_{(5)}$ and for $p>5$ we have $\mathbb{Z}[1/30]\subseteq\mathbb{Z}_{(p)}$. In light of Lemma 8, the $3\times 3$ projection matrices over $\mathbb{Z}[1/462]$ are uncolorable by Theorem 3, while those over $\mathbb{Z}[1/30]$ are uncolorable by [Bub:Schutte]. In either case, we can embed an uncolorable set in $\operatorname{Proj}(\mathbb{M}_{3}(\mathbb{Z}_{(p)}))$, and the image of this set remains uncolorable over $\mathbb{Z}_{p}$ and $\mathbb{F}_{p}$.
@@
However, our computational exploration (Table 1) has failed to produce an uncolorable set in $\mathcal{S}_{3}(6)$ to very large norm, suggesting that perhaps this set has a KS coloring. If $\mathcal{S}_{3}(6)$ turns out to be KS colorable, it is far from clear what property would separate those values of $N$ that produce KS colorable sets from those that do not. For instance, it was shown by Salt [Salt, Table 3.2] that $\mathcal{S}_{3}(714)$ is also KS uncolorable, where $714=2\cdot 3\cdot 7\cdot 17$.